# MRD: Using Physically Based Differentiable Rendering to Probe Vision Models for 3D Scene Understanding

### v3

@@ 1.1. Related work
Synthesis-based explanation methods have a long history in vision science and computer vision. The central idea is to understand the representations of image-computable vision models by synthesizing new images via an optimization process that minimizes some objective function. (Other visualization-based explanation methods primarily highlight pixels in the target image that are somehow correlated with a model’s decision (Simonyan et al., 2014; Selvaraju et al., 2017; Ribeiro et al., 2016), but will not be discussed further here). In many previous applications, this involved pixel-based gradient ascent: starting from white-noise images, and iteratively adjusting the pixels. This approach has been used to generate equivalence classes (metamers) for image-computable vision models, leading to significant impact in both computer vision and human vision science (Balas et al., 2009; Portilla and Simoncelli, [n. d.]; Freeman and Simoncelli, 2011; Wallis et al., 2016; Boehm et al., 2025; Broderick et al., 2023; Walton et al., 2021; Balas, 2006; Rosenholtz et al., 2012; Feather et al., 2023, 2019). It is also possible to synthesize images that efficiently discriminate between competing models (Golan et al., 2020; Wang and Simoncelli, 2008; Berardino et al., 2017; Feather et al., 2025; Boehm et al., 2025). Feature visualization and related techniques (Erhan et al., 2009; Mahendran and Vedaldi, 2015; Olah et al., 2017; Nguyen et al., 2019) are also forms of synthesis-based explanation. Here, images are typically generated to maximize the activation of a network layer or node. These methods are argued to provide qualitative insights into network activations and therefore explanations for their decisions, though the usefulness of these methods in providing humans with unique explanatory information has recently been called into question (Zimmermann et al., 2021; Borowski et al., 2021; Geirhos et al., 2024). Another approach has been proposed by Veeravasarapu et al. (Veeravasarapu et al., 2017) by training a GAN using a Bayesian generative model receiving a set of scene parameters and a CNN as a discrimnator to optimize the generative model via Bayesian updates to produce 3D scenes in a probabilistic manner.
@@
Physically based rendering is an area of computer graphics that seeks to solve physically based light transport often including light-surface or light-volume interactions formulated by the rendering equation ((Kajiya, 1986); see Methods). The rendering equation is a recursive formulation, which results in a high-dimensional integration problem with no analytic solution. Approximating the equation requires Monte Carlo methods alongside physical models accounting for energy conservation, i.e., the strength of the light falls off the farther it travels and the number of times it reflects or refracts from a surface. Differentiable rendering allows the optimization of the physical properties of a scene by differentiating the Monte Carlo integration in the rendering equation. The first approaches differentiating the rasterization pipeline appeared in 2014 (Loper and Black, 2014), allowing the reconstruction of scene parameters such as an object’s shape and BSDF. Third-generation differentiable renderers include Mitsuba3 (Jakob et al., 2022), PyTorch3D (Ravi et al., 2020), nvdiffrast (Laine et al., 2020), and psdr (Zhang et al., 2020). However, unlike PyTorch3D and nvdiffrast, Mitsuba3 and psdr can differentiate through integrators solving physically based rendering (path tracing) while preserving derivatives of the operations, allowing the use of gradient-based optimization methods in PBDR. Actually getting general physically based differentiable rendering to work has a number of technical challenges, caused, for example, by discontinuities with changes in lighting or due to edges, which can interfere with or bias gradient computation. Mitigating these discontinuities and making PBDR computationally tractable are active research areas in the field of differentiable rendering (Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024; Jakob, [n. d.]); we build upon these advances here.
@@ 2.1. Physically Based Differentiable Rendering
where $X_{i}$ are samples drawn from a distribution with density pp. Variance reduction strategies such as multiple importance sampling and stratified sampling improve estimator robustness (Veach, [n. d.]).
@@
A more general formulation of light transport is obtained by rewriting the RE in the space of all possible light paths $\Omega$ rather than per-surface interactions. Following Veach (Veach, [n. d.]), the contribution of pixel $j$ can be written as
@@
The derivative includes contributions from (1) the BRDF parameters, (2) geometric changes affecting $\mathbf{x}$, normal directions, and thus $\cos\theta_{i}$, (3) derivatives of incident radiance $L_{i}$, and (4) visibility terms, which introduce discontinuities that must be handled via techniques such as soft visibility, edge sampling, or regularization (Jakob et al., 2022; Vicini et al., 2021). Equation (8) forms the foundation of physically based differentiable rendering. It enables gradient-based optimization over complex, physically grounded image formation models and underlies the inverse-rendering procedure used in our experiments.
@@ 2.3. Scene
Because we are using Mitsuba (Jakob et al., 2022)(see below), we need to define our 3D scene in a Python dictionary or use the XML scene description. In contrast, neural networks do not require an initial scene configuration; we can learn the scene’s assets directly from the images. Therefore, with Mitsuba, we need to initialize a scene that we adjust with gradient updates. In this section, we will introduce the assets and the scene setup.
@@
The scene for the material reconstruction differs from the shape reconstruction. We place an object on a checkerboard and light the scene with a single environment map. The object will start from an initial, textured Principled BSDF (Disney BSDF) (Burley, [n. d.], 2015).
@@ 2.4. Implementation Details
We implement all of our experiments using Python 3.12, PyTorch (Paszke et al., 2019), and Mitsuba 3 (Jakob et al., 2022), which is a suitable PBDR for our experiments, and mitigates many of the discontinuity issues mentioned above ((Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024)). For all of our experiments, we also define an early stopping criterion: the experiment stops if the loss does not improve for 50 epochs. We run all of our shape reconstruction experiments on a workstation with a single RTX 4090 with 24 GB of VRAM using Ubuntu 22.04.5 LTS. For our material reconstruction experiments, more VRAM is required due to the path tracing integrator; for these experiments, we resort to an RTX 6000 with 48 GB of VRAM running on an identical system environment. We use pretrained models and implementations of these if available (Geirhos et al., 2019; Radford et al., 2021). Our code will be available here: https://github.com/ag-perception-wallis-lab/MRD.
@@ 2.5. Baseline experiments
The baselines use pixel-based losses. For shape, that is mean-absolute error, which works overall because it is more robust to outliers. For translucent material reconstruction, we have to use the path tracing integrator within Mitsuba (Jakob et al., 2022). This integrator has a substantially larger memory footprint than the other integrators present in Mitsuba. However, recent integrators do not allow differentiation of parameters that affect transmittance, therefore translucency. Deng et al. introduce the Dual-Buffer loss (Deng et al., 2022), which exhibits good convergence properties while requiring fewer samples per pixel, making path tracing’s usage possible. The Dual-Buffer loss computes the loss between two rendered images $\hat{y},\hat{y}\prime$ and the ground truth view $y$,
@@ 3.1. Methods
The Principled BSDF (Burley, [n. d.], 2015) was introduced to provide a production-friendly, art-directable surface model that remains physically plausible while exposing an intuitive parameter set. Rather than representing a single physical model, the Principled BSDF blends a small set of carefully chosen analytic lobes – diffuse, retro-reflective, specular microfacet, clearcoat, sheen, and transmission – into a unified framework with consistent energy conservation. Each lobe is designed to behave reasonably across the full parameter domain, enabling smooth transitions between dielectric and metallic appearances via a single metalness control while preserving reciprocity and approximate Fresnel behavior (which also benefits the optimization due to predictable changes).
@@ Acknowledgments
We thank Joshua Martin for feedback on the manuscript. We further thank the following artists and authors for providing the assets for non-commercial purposes: Delatronic (Material Dragon), Stanford’s Dragon model (Shape Dragon), Rigsters’ Lion crushing a serpent (Lion Statue), and Blender’s Suzanne. For the Hallstatt envmap, we thank Bernhard Vogl. The other three environment maps were available on HDRI Haven. The Aurora and Brushed Metal materials were measured and provided by Dupey and Jakob (Dupuy and Jakob, 2018).

### v4

@@ 1.1. Related work
Synthesis-based explanation methods have a long history in vision science and computer vision. The central idea is to understand the representations of image-computable vision models by synthesizing new images via an optimization process that minimizes some objective function. (Other visualization-based explanation methods primarily highlight pixels in the target image that are somehow correlated with a model’s decision (Simonyan et al., 2014; Selvaraju et al., 2017; Ribeiro et al., 2016), but will not be discussed further here). In many previous applications, this involved pixel-based gradient ascent: starting from white-noise images, and iteratively adjusting the pixels. This approach has been used to generate equivalence classes (metamers) for image-computable vision models, leading to significant impact in both computer vision and human vision science (Balas et al., 2009; Portilla and Simoncelli, 2000; Freeman and Simoncelli, 2011; Wallis et al., 2016; Boehm et al., 2025; Broderick et al., 2023; Walton et al., 2021; Balas, 2006; Rosenholtz et al., 2012; Feather et al., 2023, 2019). It is also possible to synthesize images that efficiently discriminate between competing models (Golan et al., 2020; Wang and Simoncelli, 2008; Berardino et al., 2017; Feather et al., 2025; Boehm et al., 2025). Feature visualization and related techniques (Erhan et al., 2009; Mahendran and Vedaldi, 2015; Olah et al., 2017; Nguyen et al., 2019) are also forms of synthesis-based explanation. Here, images are typically generated to maximize the activation of a network layer or node. These methods are argued to provide qualitative insights into network activations and therefore explanations for their decisions, though the usefulness of these methods in providing humans with unique explanatory information has recently been called into question (Zimmermann et al., 2021; Borowski et al., 2021; Geirhos et al., 2024). Another approach has been proposed by Veeravasarapu et al. (Veeravasarapu et al., 2017) by training a GAN using a Bayesian generative model receiving a set of scene parameters and a CNN as a discrimnator to optimize the generative model via Bayesian updates to produce 3D scenes in a probabilistic manner.
@@
Physically based rendering is an area of computer graphics that seeks to solve physically based light transport often including light-surface or light-volume interactions formulated by the rendering equation ((Kajiya, 1986); see Methods). The rendering equation is a recursive formulation, which results in a high-dimensional integration problem with no analytic solution. Approximating the equation requires Monte Carlo methods alongside physical models accounting for energy conservation, i.e., the strength of the light falls off the farther it travels and the number of times it reflects or refracts from a surface. Differentiable rendering allows the optimization of the physical properties of a scene by differentiating the Monte Carlo integration in the rendering equation. The first approaches differentiating the rasterization pipeline appeared in 2014 (Loper and Black, 2014), allowing the reconstruction of scene parameters such as an object’s shape and BSDF. Third-generation differentiable renderers include Mitsuba3 (Jakob et al., 2022a), PyTorch3D (Ravi et al., 2020), nvdiffrast (Laine et al., 2020), and psdr (Zhang et al., 2020). However, unlike PyTorch3D and nvdiffrast, Mitsuba3 and psdr can differentiate through integrators solving physically based rendering (path tracing) while preserving derivatives of the operations, allowing the use of gradient-based optimization methods in PBDR. Actually getting general physically based differentiable rendering to work has a number of technical challenges, caused, for example, by discontinuities with changes in lighting or due to edges, which can interfere with or bias gradient computation. Mitigating these discontinuities and making PBDR computationally tractable are active research areas in the field of differentiable rendering (Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024; Jakob et al., 2022b); we build upon these advances here.
@@ 2.1. Physically Based Differentiable Rendering
where $X_{i}$ are samples drawn from a distribution with density pp. Variance reduction strategies such as multiple importance sampling and stratified sampling improve estimator robustness (Veach, 1998).
@@
A more general formulation of light transport is obtained by rewriting the RE in the space of all possible light paths $\Omega$ rather than per-surface interactions. Following Veach (Veach, 1998), the contribution of pixel $j$ can be written as
@@
The derivative includes contributions from (1) the BRDF parameters, (2) geometric changes affecting $\mathbf{x}$, normal directions, and thus $\cos\theta_{i}$, (3) derivatives of incident radiance $L_{i}$, and (4) visibility terms, which introduce discontinuities that must be handled via techniques such as soft visibility, edge sampling, or regularization (Jakob et al., 2022a; Vicini et al., 2021). Equation (8) forms the foundation of physically based differentiable rendering. It enables gradient-based optimization over complex, physically grounded image formation models and underlies the inverse-rendering procedure used in our experiments.
@@ 2.3. Scene
Because we are using Mitsuba (Jakob et al., 2022a)(see below), we need to define our 3D scene in a Python dictionary or use the XML scene description. In contrast, neural networks do not require an initial scene configuration; we can learn the scene’s assets directly from the images. Therefore, with Mitsuba, we need to initialize a scene that we adjust with gradient updates. In this section, we will introduce the assets and the scene setup.
@@
The scene for the material reconstruction differs from the shape reconstruction. We place an object on a checkerboard and light the scene with a single environment map. The object will start from an initial, textured Principled BSDF (Disney BSDF) (Burley and Studios, 2012; Burley, 2015).
@@ 2.4. Implementation Details
We implement all of our experiments using Python 3.12, PyTorch (Paszke et al., 2019), and Mitsuba 3 (Jakob et al., 2022a), which is a suitable PBDR for our experiments, and mitigates many of the discontinuity issues mentioned above ((Nimier-David et al., 2020; Vicini et al., 2021; Nicolet et al., 2021; Zhang et al., 2023, 2024)). For all of our experiments, we also define an early stopping criterion: the experiment stops if the loss does not improve for 50 epochs. We run all of our shape reconstruction experiments on a workstation with a single RTX 4090 with 24 GB of VRAM using Ubuntu 22.04.5 LTS. For our material reconstruction experiments, more VRAM is required due to the path tracing integrator; for these experiments, we resort to an RTX 6000 with 48 GB of VRAM running on an identical system environment. We use pretrained models and implementations of these if available (Geirhos et al., 2019; Radford et al., 2021). Our code will be available here: https://github.com/ag-perception-wallis-lab/MRD.
@@ 2.5. Baseline experiments
The baselines use pixel-based losses. For shape, that is mean-absolute error, which works overall because it is more robust to outliers. For translucent material reconstruction, we have to use the path tracing integrator within Mitsuba (Jakob et al., 2022a). This integrator has a substantially larger memory footprint than the other integrators present in Mitsuba. However, recent integrators do not allow differentiation of parameters that affect transmittance, therefore translucency. Deng et al. introduce the Dual-Buffer loss (Deng et al., 2022), which exhibits good convergence properties while requiring fewer samples per pixel, making path tracing’s usage possible. The Dual-Buffer loss computes the loss between two rendered images $\hat{y},\hat{y}\prime$ and the ground truth view $y$,
@@ 3.1. Methods
The Principled BSDF (Burley and Studios, 2012; Burley, 2015) was introduced to provide a production-friendly, art-directable surface model that remains physically plausible while exposing an intuitive parameter set. Rather than representing a single physical model, the Principled BSDF blends a small set of carefully chosen analytic lobes – diffuse, retro-reflective, specular microfacet, clearcoat, sheen, and transmission – into a unified framework with consistent energy conservation. Each lobe is designed to behave reasonably across the full parameter domain, enabling smooth transitions between dielectric and metallic appearances via a single metalness control while preserving reciprocity and approximate Fresnel behavior (which also benefits the optimization due to predictable changes).
@@ Acknowledgments
We thank Joshua Martin for feedback on the manuscript. We further thank the following artists and authors for providing the assets for non-commercial purposes: Delatronic (Material Dragon), Stanford’s Dragon model (Shape Dragon), Rigsters’ Lion crushing a serpent (Lion Statue), and Blender’s Suzanne. For the Hallstatt envmap, we thank Bernhard Vogl. The other three environment maps were available on HDRI Haven. The Aurora and Brushed Metal materials were measured and provided by Dupuy and Jakob (Dupuy and Jakob, 2018).